Platform EPC:
- Lights above the platform will be installed asap (Jerome Cantet, Andy Hamy EN/EL) - Quote accepte, budget allocated. Timeline Beginning of August.
- Cable team is back in the hall: Remaining cable work (extensions and new connectors total 100 cables) will finish 20th July -> start of HWC by OP
- Nicolas retarted communication last week after the lock-out for the endoscopie intervention on the elements and shielding convined within the shielding walls between the HRS and GPS sep zones and target area (Stefano Marzari).
- Modification ex-false EPC floor in 170 still to be seen with Severin (Erwin)
197 Power supplies consolidation:
- Renforcement of the metal floor structure done and all ready to remove the old power supplies. EPC has started and is in contact with tranport for the handling of the existing racks (Antony Dillenschneider, Christophe Mutin EPC, Yann Seraphin THE)
- Modification of the floor asap after removal existing racks: Severin - news on the timeline?
- Old MISTRAL power supply now fully disconnected from AC, DC and cooling water. Can be removed (plans?) (Hanne Heylen / Dave Lunney) - attention: needs to be done in the coming weeks, before or in parallel with the removal of the other racks
REX/HIE ISOLDE:
- In the frame of the REX LLRF consolidation the REX LINAC RF cavities chacterization measurements were carried out last week by Daniel Valuch and Giampaolo Piccinini. All necessary data taken.
- Actions on the REX/HIE list:
- HIE: replacement prim. pumps CMs, backing & roughing lines - Oct/Nov (after LHC LS high peak work).
- Started: Modification of the chamber of the XL9GP Quad (last in REX LINAC before CM1/DBox) (Nicolas Thaus TE/VSC). The chamber has been taken out and will be modified with new full metal flanges. The Quad itself will also be taken out (date to be defined) for MSC (Antony Newborough) to refurbish it and modify at tthe same time the protective covers, etc. The vacuum chamber will be mounted back into the Quad when out of the machine (welding end flange)
- Additional NEG pump is to be added onto the XLN2 DBox: After re-installation end of '27.
- Removal of the obsolete 9GAP 202MHZ amplifier workaround cables (2015) from the mezzanine/MR-Tof area into the tunnel to the 9GAP structure (~11x) in the cause of this week (in parallel with RILIS cable team work). For this for a few hours the trench next to RC6 going into the HIE tunnel will be opened. Exact day and hr will be communicated asap.
- REX Low E development:
- Replacement of the leaking REX mass slits actuator for a bellow type one - decision to be taken on the type (Fredrik in strong favour of the manual version). HAs this been subject of the first SGUI meeting on Wedn 08/07 and what is the outcome?
- RILIS cable tray and cables rerouting activities: Scaffolding + garde corps have been installed. Work on heights this Thursday (no crane). In close collaboartion with the WISARD experiment.
- For installation of the CHROME related cables a scaffodling has been put up along the VITO and over the CAO line around the REXTRAP cage. Work will start with the availability of cables teams this week (after/with RILIS) and last ~2 weeks (Julien Allamandy teams).
- XT01 MINIBALL AC power distribution intervention last week done. Brought up to standard of a permanent installation (budget allocated).
- BI workzone, for HIE (and other) instrumentation, workspace is allocated between the XT03 and XT02 beamlines: 2 work benches and test set-up have been installed. Network connections been allocated (J. Tassan-Viol)
- attention: In the frame of the CHROME cable installation work the workbench next to the XT03 has been moved for a few days this week to give the cable team access to the positions of the 2 new monitors along the XT03 line. BI colleagues have been informed. Same for the area between XT01 and XT02: The wooden BI crate has been moved to the XT01 side to give access to the XT02 monitors installation.
Separator zones HRS & GPS:
- GPS concrete dust and debris cleaned up (from core drilling)
Fiber Optics work:
- 2nd phase of the White Rabbit fiber optics installation (Nenad Simic) has been scheduled for wk34 with a VIC on the 12th Aug.
Fume cupboard exhaust GLM/GHM zone:
- Options have been discussed and agreed on between CV (Nina), Transport (Yann/Filippo) and ISOLDE Tech coordination.
- Timeline: in the cause of '26